# RHuman: Real-Time 3D Human Appearance Rendering from a Single Image

### v2

@@ Abstract
Rendering 3D human appearance in different views is crucial for achieving holographic communication and immersive VR/AR. Existing methods either rely on multi-camera setups or have low-quality rendered images from a single image. In this paper, we propose RHuman, the first approach for real-time inference and rendering of photorealistic 3D human appearance from a single image. The core of our approach is to combine the strengths of implicit texture fields and explicit neural rendering with our novel representation, namely Z-map. Based on this, we present an end-to-end network that performs high-fidelity color reconstruction of visible areas and provides reliable color inference for occluded regions. To further enhance the 3D perception ability of our network, we leverage the Fourier occupancy field as a prior for generating the texture field and providing a sampling surface in the rendering stage. We also propose a consistency loss and a spatio-temporal fusion strategy to ensure the multi-view coherence. Experimental results show that our method outperforms the state-of-the-art methods on both synthetic data and challenging real-world images, in real time. The source code will be available for research purposes.
@@
Given a single RGB image, RHuman can generate photorealistic 3D human appearance in real-time. We utilized the input (a) to render close views (b) and zoomed in (c). The results of pitch angle changes are shown in (d-e), and (f-h) demonstrate the outcomes of divergent views.
@@ Introduction
The rapid evolution of digital technology has notably impacted computer graphics and computer vision, particularly in 3D human appearance reconstruction. This advancement not only enhances visual experiences but also paves the way for practical applications in AR/VR, enabling interactive platforms like holographic communication and immersive social experiences. However, existing methods heavily rely on multi-camera setups [34, 19, 15, 20] or are constrained to offline operations [1, 7, 16], leaving real-time 3D human rendering from a single image as an unresolved challenge. In this paper, we aim to render the photorealistic 3D human appearance in real-time from a single image, as shown in Fig.RHuman: Real-Time 3D Human Appearance Rendering from a Single Image.
@@
In this paper, we introduce RHuman, a real-time framework for rendering human appearance from a single image, which uniquely combines flow-based rendering techniques with an implicit 3D human geometry representation to synthesize novel view images. Our method only requires a single image as input and achieves both high-quality and real-time rendering performance. Moreover, our method can generate novel view renderings of humans in various clothing without individual-specific training. Unlike previous flow-based rendering methods that omit points occluded in the source view, we introduce an intermediate representation, namely the Z-map, during the rendering process, which collects the source view depth of the rendered points of the target view and forms them to a 2D map. It helps lift the 2D image feature into a 3D texture feature field, while maintaining compatibility with 2D neural rendering networks. This unique capability enables our network to learn data-driven 2D appearance knowledge of clothed human images, resulting in more accurate renderings. By preserving the features of occluded points and leveraging the Z-map, our method effectively resolves depth ambiguities. Additionally, we employ an efficient 3D object representation known as Fourier occupancy fields (FOF) [4], which explicitly represents a 3D object as a multi-channel image. It can serve both as a prior for texture field generation and as a sampling surface during the rendering stage, avoiding the high computational costs caused by dense sampling. In order to ensure the consistency between multiple views and reduce the jittering phenomenon between neighboring frames, we propose a consistency loss to regularize this process and design a spatio-temporal fusion strategy to enhance this in practical applications. Our method paves the way for the practical applications of AR/VR, which can be applied in holographic communication in the future. Source code will be available for research purposes.
@@
We present a novel system for high-quality, real-time synthesis of human novel view images with only a single RGB image input. To the best of our knowledge, this is the first system to restore the full-body appearance of a 3D human in real time from a single image, paving the way for practical applications in AR/VR.
We propose RHuman, an end-to-end CNN-based neural rendering method that combines the strengths of implicit texture field and explicit neural rendering, which can produce results with both 3D consistency and high visual quality.
@@
We propose a consistency loss to ensure the multi-view coherence and reduce the jittering phenomenon, and design a spatio-temporal fusion strategy to enhance this in practical applications.
@@ 1.1 Monocular 3D Human Reconstruction
Human reconstruction has long been a concern in the domain of computer vision and graphics. Methods for predicting human bodies from a single image represent 3D shapes primarily by estimating the properties of points in 3D space. Some previous works [33, 22] directly predict the occupancy field of a given segment in space through regression networks, but such methods have high memory requirements, which limits the spatial resolution of shape estimation. Another type of methods involves using implicit function networks to remove resolution constraints. Saito et al. [16] proposed an implicit function based on pixel-aligned features to reconstruct a 3D human from a single image. PIFuHD [17] further introduces normal maps to improve geometric details. Subsequent methods [32, 27, 26, 30] improve the robustness of the results by incorporating parametric SMPL [12] priors. Although these methods can achieve realistic results, they often require significant computational resources and are time-consuming. Feng et al. [4, 5] proposed an efficient 3D geometric representation called Fourier Occupancy Field, which can represent a high-quality 3D geometry as a multi-channel image. It establishes a strong link between 3D geometry and 2D images, significantly reducing the computational requirements for reconstruction. However, one limitation is its inability to estimate texture during the geometry reconstruction process.
@@ 1.2 Human Novel View Synthesis
In this paper, we contribute the first real-time and high-quality human novel view synthesis system from a single RGB image, which combines the strengths of implicit texture field and explicit neural rendering, opening up novel possibilities for holographic communication and immersive social interaction. Tab. 1 presents a comparative analysis focusing on key features: support for monocular input, real-time processing capabilities, full-body rendering, and high-fidelity novel view synthesis. As indicated in the table, our method is unique in its support for all these features, distinguishing it from other approaches.
2 Preliminary
FOF [4] introduces a novel, powerful, efficient and flexible 3D geometry representation called Fourier occupancy field, for real-time and accurate human reconstruction from a single image. The Fourier Occupancy Field represents a 3D object with a 2D field orthogonal to the view direction. Specifically, FOF approximates the 1D occupancy signal of point on the 2D plane orthogonal to the direction of the view with a subspace spanned by the first basis functions of a convergent Fourier series:
where are coefficients of basis functions and , respectively. Then, the human geometry is represented as
where is the vector of the first basis functions spanning the approximation subspace, and is the Fourier coefficient vectors for the 1D occupancy signal at . FOF stores a 3D object as a multi-channel image that can be quickly restored to a 3D occupied field by matrix multiplication, which facilitates prediction using a CNN.
@@ 3.2.2 Novel View Rendering
3.2.3 Spatio-Temporal Fusion Strategy
To maintain consistency during the rotation of the view, we introduce a spatio-temporal fusion strategy for generating the free-view video stream. Specifically, during the rotation process, we keep the image of the previous frame , and use Eq. 11 to get the surface points that are visible in both the current and previous frames. Subsequently, the color of in the current frame image is interpolated as:
@@ 3.3 Training
We use synthetic data to train RHuman during the training stage. As illustrated in Fig. 1, an input image along with its corresponding FOF and normal map, sampled from the training set, serves as the input. The encoder then generates the corresponding texture field features, denoted as . With the camera parameters (comprising camera direction and position), we compute the warping flow which contains the novel view information we need, enabling the generation of an image of any view with the decoder. In each gradient step, we synthesize two images , with the same texture field features from a reference camera and an additional camera to perform multi-view consistency supervision.
@@ 4 Experiments
4.1 Real-time Rendering System
Our pipeline, designed for real-time monocular human novel view synthesis, operates as follows: Images are captured from the video stream using OpenCV [3] and preprocessed to a resolution of . We then apply RVM [11] for human mask extraction, effectively removing the background, and use HMR2.0 [8] for SMPL model estimation. Subsequently, the FOF-SMPL [4] is utilized to infer the occupation field. CUDA is employed to accelerate the rendering of depth, normal, and IUV maps. The final high-fidelity novel view images are rendered in real-time with the camera parameters and the previously processed data. Our pipeline, implemented with TensorRT on a single RTX-3090 GPU, achieves a performance of 24+ FPS, with the potential for further enhancement via additional GPUs. Fig. 2 showcase some of the real-time reconstruction results. Although our method trains to process each frame separately using synthetic data, it can still provide reasonable temporal consistency.
| Component | Streamer | Segmentation | FOF-SMPL | Additional | Inference(Ours) |
| & SMPL | information | ||||
| Time(ms) | 4.21 | 10.59 | 8.77 | 0.93 | 13.13 |
Tab. 2 details the execution time for each component of our pipeline. Currently, we implement our pipeline serially, achieving an overall runtime of approximately 37ms. We have combined the neural network and recovering of the occupancy field into a single component, optimized with TensorRT for enhanced acceleration. Moreover, there is potential for substantially increasing the pipeline’s processing speed through parallel implementation of each component.
@@ 4.2 Comparisons
Qualitative results. Fig. 3 and Fig. 4 presents qualitative comparisons of our approach to the baseline on the 2k2k dataset [6] and the THuman2.0 dataset [29], respectively. In addition, we tested the models on in-the-wild images, as shown in Fig. 5. PIFu demonstrates the ability to estimate reasonably accurate colors, but its performance is hindered by limitations in geometric estimation, resulting in suboptimal image quality. On the other hand, while SHERF can produce results with accurate pose, it encounters challenges in generating desirable outcomes for individuals wearing loose clothing, mainly due to its heavy reliance on SMPL priors. Additionally, since SHERF estimates colors for sampling points independently, it does not effectively consider the interrelationship between these points, often leading to ambiguous results. Fig. 6 shows a qualitative comparison with traditional texture mapping methods. 3DTexture has a noticeable concatenation artifacts in the multi-view settings and cannot render occluded areas in the input view. Thanks to Z-map, Rhuman is able to incorporate geometric priors to enhance the rendering of occluded areas by 3D information. Furthermore, with the aid of neural rendering, Rhuman can effectively compensate for any inaccuracies in the geometric estimates. As a result, Rhuman is capable of synthesizing high-fidelity results that exhibit remarkable 3D consistency.
Quantitative evaluations. We selected 105 models in the 2k2k and THuman2.0 test sets for evaluation, respectively. Tab. 3 shows the performance of the method when the novel view is in close proximity to the source view (with a difference of 45°), and Tab. 4 shows the performance when the novel view is significantly different from the source view (with a difference of 90°). It is evident that RHuman outperforms SHERF and PIFu across all evaluation metrics, regardless of whether there are minor or major changes in the viewing angle.
@@ 4.3 Ablation study
| SSIM | PSNR | LPIPS | |
@@
Effects of consistency strategy. We add a consistency loss and a spatio-temporal fusion strategy to ensure the multi-view consistency of the results. To explore the impact of the spatio-temporal fusion strategy on the results, we retrain a variant without the consistency loss and use the method described in [23] to draw EPI images to qualitatively evaluate the view consistency. Fig. 9 compares the results with and without the spatio-temporal fusion strategy, and it can be seen that our spatio-temporal fusion strategy effectively enhances the view consistency during rotation.
4.4 Results on 1K Resolution
Our method in the paper was trained and tested on a resolution of . We also explored the performance of the method at a higher resolution of 1K 1K. Specifically, we created a 1K dataset and trained our model on it. Fig. 12 presents the results of our method at this higher resolution. It can be seen that the use of high-resolution images significantly improves the sharpness of the results and enhances visual effects.
4.5 Failure Cases
Although our method makes full use of the 3D geometry information in the reconstructed data, there may still be artifacts caused by inaccurate geometry estimation. In addition, due to the absence of an explicit segmentation of skin parts and clothing parts, there are cases where the network may produce inaccurate color estimations on specific body parts. Fig. 13 illustrates some examples of these failure cases. In the future, we plan to explore additional supervision methods to enhance the network’s ability to recognize human body parts more effectively and thereby improve the quality of the output images.
@@ 5 Conclusion and Discussion
Conclusion. Rendering 3D human appearance from a single image in real time is important for achieving holographic communication and immersive VR/AR. We propose a novel method, combining the strengths of implicit texture field and explicit neural rendering, for real-time inference and rendering of realistic 3D human appearance from a monocular RGB image. We propose a new representation, Z-map, to alleviate depth ambiguities in rendering and enable high-fidelity color reconstruction. We also design a consistency loss and a spatio-temporal fusion strategy to ensure the multi-view coherence and reduce the jittering phenomenon. Experimental results show that the proposed method achieves state-of-the-art performance on both synthetic data and real-world images.

### v4

@@ Abstract
Rendering 3D human appearance from a single image in real-time is crucial for achieving holographic communication and immersive VR/AR. Existing methods either rely on multi-camera setups or are constrained to offline operations. In this paper, we propose R2Human, the first approach for real-time inference and rendering of photorealistic 3D human appearance from a single image. The core of our approach is to combine the strengths of implicit texture fields and explicit neural rendering with our novel representation, namely Z-map. Based on this, we present an end-to-end network that performs high-fidelity color reconstruction of visible areas and provides reliable color inference for occluded regions. To further enhance the 3D perception ability of our network, we leverage the Fourier occupancy field as a prior for generating the texture field and providing a sampling surface in the rendering stage. We also propose a consistency loss and a spatial fusion strategy to ensure the multi-view coherence. Experimental results show that our method outperforms the state-of-the-art methods on both synthetic data and challenging real-world images, in real-time. The project page can be found at http://cic.tju.edu.cn/faculty/likun/projects/R2Human.
@@
Given a single RGB image, R2Human can generate photorealistic 3D human appearance in real-time. We utilized the input (a) to render close views (b) and zoomed in (c). The results of pitch angle changes are shown in (d-e), and (f-h) demonstrate the outcomes of divergent views.
@@ Introduction
Rendering 3D human appearance from a single image in real-time not only enhances visual experiences but also paves the way for practical applications in AR/VR, enabling users to engage in immersive experiences with timely feedback. However, existing methods heavily rely on multi-camera setups [40, 23, 19, 24] or are constrained to offline operations [2, 9, 20, 1], leaving real-time 3D human rendering from a single image as an unresolved challenge. In this paper, we aim to translate a 2D human image into a vibrant 3D appearance in real-time, facilitating holographic communication, and enhancing user experience and immersion in VR/AR.
@@
In this paper, we introduce R2Human, a real-time framework for rendering human appearance from a single image, which uniquely combines flow-based rendering techniques with an implicit 3D human geometry representation to synthesize novel view images. Our method only requires a single image as input and achieves both high-quality rendering performance, as shown in Fig.R2Human: Real-Time 3D Human Appearance Rendering from a Single Image. Unlike previous human avatar animation typically involves creating actions beyond real-world scenarios, our focus lies in enhancing existing visual experiences by transforming real 2D human images into immersive 3D appearances in real-time. By converting existing 2D images to 3D, we cleverly avoids the increase in computational consumption caused by animation rendering and the decrease in the authenticity of the results. Moreover, our method can generate novel view renderings of humans in various clothing without individual-specific training. We introduce an intermediate representation, namely the Z-map, during the rendering process, which collects the source view depth of the rendered points of the target view and forms them to a 2D map. It helps lift the 2D image feature into a 3D texture feature field, while maintaining compatibility with 2D neural rendering networks. This unique capability enables our network to learn data-driven 2D appearance knowledge of clothed human images, resulting in more accurate renderings. By preserving the features of occluded points and leveraging the Z-map, our method effectively resolves depth ambiguities. Additionally, we employ an efficient 3D object representation known as Fourier occupancy fields (FOF) [5], which explicitly represents a 3D object as a multi-channel image. It can serve both as a prior for texture field generation and as a sampling surface during the rendering stage, avoiding the high computational costs caused by dense sampling. In order to ensure the consistency between multiple views and reduce the jittering phenomenon between neighboring views, we propose a consistency loss to regularize this process and design a spatial fusion strategy to enhance this in practical applications. Our method paves the way for the practical applications of AR/VR, which can be applied in holographic communication in the future. Source code will be available for research purposes.
@@
We present a novel system for high-quality, real-time synthesis of human novel view images with only a single RGB image input. To the best of our knowledge, this is the first system to restore the full-body appearance of a 3D human in real-time from a single image, paving the way for practical applications in AR/VR.
We propose R2Human, an end-to-end CNN-based neural rendering method that combines the strengths of implicit texture field and explicit neural rendering, which can produce results with both 3D consistency and high visual quality.
@@
We propose a consistency loss to ensure the multi-view coherence and reduce the jittering phenomenon, and design a spatial fusion strategy to enhance this in practical applications.
@@ 1.1 Monocular 3D Human Reconstruction
Human reconstruction has long been a concern in the domain of computer vision and graphics. Methods for predicting human bodies from a single image represent 3D shapes primarily by estimating the properties of points in 3D space. Some previous works [39, 27] directly predict the occupancy field of a given segment in space through regression networks, but such methods have high memory requirements, which limits the spatial resolution of shape estimation. Another type of methods involves using implicit function networks to remove resolution constraints. Saito et al. [20] proposed an implicit function based on pixel-aligned features to reconstruct a 3D human from a single image. PIFuHD [21] further introduces normal maps to improve geometric details. Subsequent methods [38, 33, 32, 36] improve the robustness of the results by incorporating parametric SMPL [15] priors. Albahar et al. [1] combine diffusion model to obtain detailed geometry and texture, but the reconstruction effect of loose clothing is not good. Some methods [41, 7, 10, 31, 12, 11] take advantage of the features of SMPL to align different poses to obtain detailed reconstructions from video, but cannot adapt to single image settings. Although these methods can achieve realistic results, they often require significant computational resources and are time-consuming. Feng et al. [5, 6] proposed an efficient 3D geometric representation called Fourier Occupancy Field, which can establish a strong link between 3D geometry and 2D images, significantly reducing the computational requirements for reconstruction. However, one limitation is its inability to estimate texture during the geometry reconstruction process. Other work [26, 16] relies on RGBD cameras to provide depth information, thus ensuring real-time performance and reconstruction accuracy. But their results rely too much on the SMPL model, making it impossible to adapt to a variety of clothes, and the use of RGBD cameras also limits its application in daily life.
@@ 1.2 Human Novel View Synthesis
In this paper, we contribute the first real-time and high-quality human novel view synthesis system from a single RGB image, which combines the strengths of implicit texture field and explicit neural rendering and translates a 2D human image into a vibrant 3D appearance in real-time, facilitating holographic communication, and enhancing user experience and immersion in VR/AR. Tab. 1 presents a comparative analysis focusing on key features: support for monocular input, real-time processing capabilities, full-body rendering, and high-fidelity novel view synthesis. As indicated in the table, our method is unique in its support for all these features, distinguishing it from other approaches.
@@ 2.2.2 Novel View Rendering
2.2.3 Spatial Fusion Strategy
To maintain consistency during the rotation of the view, we introduce a spatial fusion strategy for generating the free-view video stream. Specifically, during the rotation process, we keep the image of the previous view , and use Eq. 9 to get the surface points that are visible in both the current and previous views. Subsequently, the color of in the current view image is interpolated as:
@@ 2.3 Training
We use synthetic data to train R2Human during the training stage. As illustrated in Fig. 1, an input image along with its corresponding FOF and normal map, sampled from the training set, serves as the input. The encoder then generates the corresponding texture field features, denoted as . With the camera parameters (comprising camera direction and position), we compute the warping flow which contains the novel view information we need, enabling the generation of an image of any view with the decoder. In each gradient step, we synthesize two images , with the same texture field features from a reference camera and an additional camera to perform multi-view consistency supervision.
@@ 3.1 Comparisons
Qualitative results. Fig. 2 and Fig. LABEL:fig:qualitative_2k2k presents qualitative comparisons of our approach to the baseline on the 2k2k dataset [8] and the THuman2.0 dataset [35], respectively. PIFu demonstrates the ability to estimate reasonably accurate colors, but its performance is hindered by limitations in geometric estimation, resulting in suboptimal image quality. On the other hand, while SHERF can produce results with accurate pose, it encounters challenges in generating desirable outcomes for individuals wearing loose clothing, mainly due to its heavy reliance on SMPL priors. Additionally, since SHERF estimates colors for sampling points independently, it does not effectively consider the interrelationship between these points, often leading to ambiguous results. Fig. 3 shows a qualitative comparison with traditional texture mapping methods. 3DTexture has a noticeable concatenation artifacts in the multi-view settings and cannot render occluded areas in the input view. Thanks to Z-map, R2human is able to incorporate geometric priors to enhance the rendering of occluded areas by 3D information. Furthermore, with the aid of neural rendering, R2human can effectively compensate for any inaccuracies in the geometric estimates. As a result, R2human is capable of synthesizing high-fidelity results that exhibit remarkable 3D consistency.
Quantitative evaluations. We selected 105 models in the 2k2k and THuman2.0 test sets for evaluation, respectively. Tab. 2 shows the performance of the method when the novel view is in close proximity to the source view (with a difference of 45°), and Tab. 3 shows the performance when the novel view is significantly different from the source view (with a difference of 90°). It is evident that R2Human outperforms SHERF and PIFu across all evaluation metrics, regardless of whether there are minor or major changes in the viewing angle.
@@ 3.2 Ablation study
| Method | SSIM | PSNR | LPIPS |
@@
Effects of consistency strategy. We add a consistency loss and a spatial fusion strategy to ensure the multi-view consistency of the results. To explore the impact of the spatial fusion strategy on the results, we retrain a variant without the consistency loss and use the method described in [28] to draw EPI images to qualitatively evaluate the view consistency. Fig. 6 compares the results with and without the spatial fusion strategy, and it can be seen that our spatial fusion strategy effectively enhances the view consistency during rotation.
3.3 Real-time Rendering System
Our pipeline, designed for real-time monocular human novel view synthesis, operates as follows: Images are captured from the video stream using OpenCV [4] and preprocessed to a resolution of . We then apply FOF-SMPL [5] to infer the occupation field. Subsequently, we employ the marching cube algorithm to extract the mesh and render the depth and normal maps. The final high-fidelity novel view images are rendered in real-time with the camera parameters and the previously processed data. Our pipeline, implemented with TensorRT on a single RTX-3090 GPU, achieves a performance of 24+ FPS, with the potential for further enhancement via additional GPUs. Fig. 9 showcase some of the real-time reconstruction results. Although our method trains to process each frame separately using synthetic data, it can still provide reasonable temporal consistency. Please refer to the demo video for more results.
3.4 Virtual Reality and Augmented Reality Applications
Our method has many meaningful applications, and we present two applications here. As shown in the top row of Fig. 10, we render the human appearance into a panoramic photo to generate the person in a virtual environment and display it in VR glasses. Specifically, we project the results into a spherical coordinate system and expand the human appearance synthesized into the panoramic image to obtain the panoramic human appearance. By rendering the 3D human appearance into the virtual environment, our method provides people with a more realistic and stunning virtual experience. In addition, our method can also be integrated into the real scene of AR applications. As shown in the bottom row of Fig. 10, given the camera parameters of each frame in the video, we can render the human appearance in a specific scene and observe the 3D human from different views. Our VR/AR examples can generate many realistic applications in certain situations, e.g., AR/VR education, immersion effects in games, virtual character communication in real scenes, etc. Also, our method of realistic 3D human appearance rendering has promising applications in immersive telepresence. In virtual conferences, the real 3D human appearance can greatly increase people’s immersion. Applying our method in these VR/AR applications is a valuable direction in the future.
@@ 4 Conclusion and Discussion
Conclusion. Rendering 3D human appearance from a single image in real-time is important for achieving holographic communication and immersive VR/AR. We propose a novel method, combining the strengths of implicit texture field and explicit neural rendering, for real-time inference and rendering of realistic 3D human appearance from a monocular RGB image. We propose a new representation, Z-map, to alleviate depth ambiguities in rendering and enable high-fidelity color reconstruction. We also design a consistency loss and a spatial fusion strategy to ensure the multi-view coherence and reduce the jittering phenomenon. Experimental results show that the proposed method achieves state-of-the-art performance on both synthetic data and real-world images.
@@
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (62122058 and 62171317), and the Science Fund for Distinguished Young Scholars of Tianjin (No. 22JCJQJC00040).References
- [1] B. AlBahar, S. Saito, H.-Y. Tseng, C. Kim, J. Kopf, and J.-B. Huang. Single-image 3d human digitization with shape-guided diffusion. In SIGGRAPH Asia 2023 Conference Papers, pp. 1–11, 2023.
- [2] T. Alldieck, M. Zanfir, and C. Sminchisescu. Photorealistic monocular 3d reconstruction of humans wearing clothing. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 1506–1515, 2022.
- [3] S. Avidan and A. Shashua. Novel view synthesis in tensor space. In Proceedings of IEEE Computer Society Conference on Computer Vision and Pattern Recognition, pp. 1034–1040. IEEE, 1997.
- [4] G. Bradski. The opencv library. Dr. Dobb’s Journal: Software Tools for the Professional Programmer, 25(11):120–123, 2000.
- [5] Q. Feng, Y. Liu, Y.-K. Lai, J. Yang, and K. Li. Fof: learning fourier occupancy field for monocular real-time human reconstruction. Advances in Neural Information Processing Systems, 35:7397–7409, 2022.
- [6] Q. Feng, Y. Liu, Y.-K. Lai, J. Yang, and K. Li. Monocular real-time human geometry reconstruction. In CAAI International Conference on Artificial Intelligence, pp. 594–598, 2022.
- [7] C. Guo, T. Jiang, X. Chen, J. Song, and O. Hilliges. Vid2avatar: 3d avatar reconstruction from videos in the wild via self-supervised scene decomposition. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 12858–12868, 2023.
- [8] S.-H. Han, M.-G. Park, J. H. Yoon, J.-M. Kang, Y.-J. Park, and H.-G. Jeon. High-fidelity 3d human digitization from single 2k resolution images. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 12869–12879, 2023.
- [9] S. Hu, F. Hong, L. Pan, H. Mei, L. Yang, and Z. Liu. Sherf: Generalizable human nerf from a single image. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pp. 9352–9364, 2023.
- [10] B. Jiang, Y. Hong, H. Bao, and J. Zhang. Selfrecon: Self reconstruction your digital avatar from monocular video. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 5605–5615, 2022.
- [11] T. Jiang, X. Chen, J. Song, and O. Hilliges. Instantavatar: Learning avatars from monocular video in 60 seconds. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 16922–16932, 2023.
- [12] W. Jiang, K. M. Yi, G. Samei, O. Tuzel, and A. Ranjan. Neuman: Neural human radiance field from a single video. In European Conference on Computer Vision, pp. 402–418. Springer, 2022.
- [13] A. Krizhevsky, I. Sutskever, and G. E. Hinton. Imagenet classification with deep convolutional neural networks. Advances in neural information processing systems, 25, 2012.
- [14] J. Lawrence, D. B. Goldman, S. Achar, G. M. Blascovich, J. G. Desloge, T. Fortes, E. M. Gomez, S. Häberling, H. Hoppe, A. Huibers, et al. Project starline: A high-fidelity telepresence system. 2021.
- [15] M. Loper, N. Mahmood, J. Romero, G. Pons-Moll, and M. J. Black. Smpl: A skinned multi-person linear model. Acm Transactions on Graphics, 34(Article 248), 2015.
- [16] Y. Lu, H. Yu, W. Ni, and L. Song. 3d real-time human reconstruction with a single rgbd camera. Applied Intelligence, 53(8):8735–8745, 2023.
- [17] B. Mildenhall, P. P. Srinivasan, M. Tancik, J. T. Barron, R. Ramamoorthi, and R. Ng. Nerf: Representing scenes as neural radiance fields for view synthesis. Communications of the ACM, 65(1):99–106, 2021.
- [18] P. Nguyen-Ha, N. Sarafianos, C. Lassner, J. Heikkilä, and T. Tung. Free-viewpoint rgb-d human performance capture and rendering. In European Conference on Computer Vision, pp. 473–491. Springer, 2022.
- [19] S. Peng, Y. Zhang, Y. Xu, Q. Wang, Q. Shuai, H. Bao, and X. Zhou. Neural body: Implicit neural representations with structured latent codes for novel view synthesis of dynamic humans. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 9054–9063, 2021.
- [20] S. Saito, Z. Huang, R. Natsume, S. Morishima, A. Kanazawa, and H. Li. Pifu: Pixel-aligned implicit function for high-resolution clothed human digitization. In Proceedings of the IEEE/CVF international conference on computer vision, pp. 2304–2314, 2019.
- [21] S. Saito, T. Simon, J. Saragih, and H. Joo. Pifuhd: Multi-level pixel-aligned implicit function for high-resolution 3d human digitization. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 84–93, 2020.
- [22] U. Sara, M. Akter, and M. S. Uddin. Image quality assessment through fsim, ssim, mse and psnr—a comparative study. Journal of Computer and Communications, 7(3):8–18, 2019.
- [23] R. Shao, L. Chen, Z. Zheng, H. Zhang, Y. Zhang, H. Huang, Y. Guo, and Y. Liu. Floren: Real-time high-quality human performance rendering via appearance flow using sparse rgb cameras. In SIGGRAPH Asia 2022 Conference Papers, pp. 1–10, 2022.
- [24] R. Shao, H. Zhang, H. Zhang, M. Chen, Y.-P. Cao, T. Yu, and Y. Liu. Doublefield: Bridging the neural surface and radiance fields for high-fidelity human reconstruction and rendering. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 15872–15882, 2022.
- [25] K. Simonyan and A. Zisserman. Very deep convolutional networks for large-scale image recognition. arXiv preprint arXiv:1409.1556, 2014.
- [26] H. Song, B. Yoon, W. Cho, and W. Woo. Rc-smpl: Real-time cumulative smpl-based avatar body generation. In 2023 IEEE International Symposium on Mixed and Augmented Reality (ISMAR), pp. 89–98. IEEE, 2023.
- [27] G. Varol, D. Ceylan, B. Russell, J. Yang, E. Yumer, I. Laptev, and C. Schmid. Bodynet: Volumetric inference of 3d human body shapes. In Proceedings of the European conference on computer vision (ECCV), pp. 20–36, 2018.
- [28] A. Vianello, J. Ackermann, M. Diebold, and B. Jähne. Robust hough transform based 3d reconstruction from circular light fields. In CVPR, 2018.
- [29] M. Waechter, N. Moehrle, and M. Goesele. Let there be color! large-scale texturing of 3d reconstructions. In Computer Vision–ECCV 2014: 13th European Conference, Zurich, Switzerland, September 6-12, 2014, Proceedings, Part V 13, pp. 836–850. Springer, 2014.
- [30] Z. Wang, A. C. Bovik, H. R. Sheikh, and E. P. Simoncelli. Image quality assessment: from error visibility to structural similarity. IEEE transactions on image processing, 13(4):600–612, 2004.
- [31] C.-Y. Weng, B. Curless, P. P. Srinivasan, J. T. Barron, and I. Kemelmacher-Shlizerman. Humannerf: Free-viewpoint rendering of moving people from monocular video. In Proceedings of the IEEE/CVF conference on computer vision and pattern Recognition, pp. 16210–16220, 2022.
- [32] Y. Xiu, J. Yang, X. Cao, D. Tzionas, and M. J. Black. Econ: Explicit clothed humans obtained from normals. arXiv preprint arXiv:2212.07422, 2022.
- [33] Y. Xiu, J. Yang, D. Tzionas, and M. J. Black. Icon: Implicit clothed humans obtained from normals. In 2022 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 13286–13296. IEEE, 2022.
- [34] X. Yang, Y. Luo, Y. Xiu, W. Wang, H. Xu, and Z. Fan. D-if: Uncertainty-aware human digitization via implicit distribution field. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pp. 9122–9132, 2023.
- [35] T. Yu, Z. Zheng, K. Guo, P. Liu, Q. Dai, and Y. Liu. Function4d: Real-time human volumetric capture from very sparse consumer rgbd sensors. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pp. 5746–5756, 2021.
- [36] M. Zhang, Q. Feng, Z. Su, C. Wen, Z. Xue, and K. Li. Joint2human: High-quality 3d human generation via compact spherical embedding of 3d joints. arXiv preprint arXiv:2312.08591, 2023.
- [37] R. Zhang, P. Isola, A. A. Efros, E. Shechtman, and O. Wang. The unreasonable effectiveness of deep features as a perceptual metric. In Proceedings of the IEEE conference on computer vision and pattern recognition, pp. 586–595, 2018.
- [38] Z. Zheng, T. Yu, Y. Liu, and Q. Dai. Pamir: Parametric model-conditioned implicit representation for image-based human reconstruction. IEEE transactions on pattern analysis and machine intelligence, 44(6):3170–3184, 2021.
- [39] Z. Zheng, T. Yu, Y. Wei, Q. Dai, and Y. Liu. Deephuman: 3d human reconstruction from a single image. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pp. 7739–7749, 2019.
- [40] T. Zhou, J. Huang, T. Yu, R. Shao, and K. Li. Hdhuman: High-quality human novel-view rendering from sparse views. IEEE Transactions on Visualization and Computer Graphics, 2023.
- [41] W. Zielonka, T. Bagautdinov, S. Saito, M. Zollhöfer, J. Thies, and J. Romero. Drivable 3d gaussian avatars. 2024.